OpenStudy (anonymous):

Auditions for a new, upcoming band in Chicago were held with 90 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 30, how many women auditioned? 50 40 60 30

OpenStudy (anonymous):

30<----- answer

OpenStudy (jackellyn):

From the word problem, create equations 90=m+w 2m-3w=30 Solve for variable m to substitute into first eqn m=90-w Substitute 2(90-w)-3w=30 Distribute 180 -2w-3w =30 Combine like terms -5w=-150 Solve for w w=-150/-5 w=30 So 30 women auditioned. @savannajadey